OpenStudy (anonymous):

For the function y = f(x), what is the ordered pair for the point on the graph when x = b - 2? (b - 2, f(b -2)) (x, f(b)) (x, b - 2) (b - 2, f(b))

myininaya (myininaya):

Your points will look like this on your curve (x,y=f(x)) You are given the x there. So just plug it in into (x,f(x)) wherever you see x put bis

OpenStudy (anonymous):

y=f(x+1) has ordered pairs of the form (x+1,y) y=f(x) has ordered pairs of the form (x,y) we know a point on y=f(x) which is (2,4) so when does (2,4)=(x+1,y) you said y=4 so we have (2,4)=(x+1,4)

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

u r welcome